# PHENIX Collaboration

### v1

@@ III.1 Particle Ratios
Enhancement of baryon production in nucleus-nucleus collisions is considered to be one of the signatures of QGP formation [14, 15, 17]. To investigate differences in baryon and meson production mechanisms, the ratios of $p/\pi$ and $K/\pi$ have been calculated. Figures 9 and 9 present comparisons of $p$/$\pi^{+}$ and $\bar{p}$/$\pi^{-}$ ratios in different centralities of large (Cu$+$Au, U$+$U) and small ($p$$+$$p$, $p$$+$Al, ${}^{3}$e$+$Au) collision systems. In central collisions of large systems $p$/$\pi$ ratios reach the values of $\approx$0.6, but in peripheral collisions the values of $p/\pi$ ratios are smaller than 0.4 in the whole $p_{T}$ range. Behavior of $p/\pi$ ratios observed in Cu$+$Au and U$+$U collision systems can be qualitatively described using recombination models [23, 24].
In small collision systems ($p$$+$Al, ${}^{3}$e$+$Au), the values of $p$/$\pi$ ratios are similar to those measured in $p$$+$$p$ collisions [43]. In ${}^{3}$e$+$Au collisions a modest centrality dependence can be seen, similar to that observed in $d$$+$Au collisions [14, 17]. The modest centrality dependence can be understood in terms of the small range of $\langle N_{\rm part}\rangle$ values relative the large range of $\langle N_{\rm part}\rangle$ values in large collision systems. The values of $p/\pi$ ratios measured in all centrality classes of $p$$+$Al collisions and in $p$$+$$p$ collisions are consistent within uncertainties. For all measurements type A uncertainties sum quadratically to type B uncertainties and are shown as a rectangles around the experimental points in Figs.1115.
@@ III.2 Nuclear-Modification Factors
Figures 13 and 13 present comparisons of identified charged-hadron $R_{AB}$ values as a function of $p_{T}$ in central and peripheral $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au, and U$+$U collisions. The $R_{AB}$ values are found to be in agreement in collisions with different geometries, but with the same $\langle N_{\rm part}\rangle$ values, indicating that identified charged-hadron production depends only on system size and not geometry.
The following features of identified charged-hadron production in $p$$+$Al collisions have been found: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than it is in ${}^{3}$e$+$Au and $d$$+$Au collisions and (ii) proton $R_{AB}$ values in $p$$+$Al collisions at the intermediate $p_{T}$ range (1.0 GeV/$c$ $<~{}p_{T}~{}<$ 2.5 GeV/$c$) are equal to unity, while in ${}^{3}$e$+$Au and $d$$+$Au collisions proton $R_{AB}$ is above unity. Differences in identified charged-hadron production between $p$$+$Al and $d/$${}^{3}$e$+$Au might be caused by the size of the $p$$+$Al system being insufficient to observe an increase in proton production.
Comparisons of identified charged-hadron $R_{AB}$ values with neutral-meson $R_{AB}$ values [47, 44, 45] in small and large collision systems are presented in Figs. 15 and 15. In large collision systems and in the ${}^{3}$e+Au collision system, proton $R_{AB}$ values are enhanced over all meson $R_{AB}$ values. The mass of the $\phi$-meson $m_{\phi}$=1019 MeV/$c^{2}$ is similar to the proton mass $m_{p}$=938 MeV/$c^{2}$, therefore the enhancement of proton $R_{AB}$ values over $\phi$-meson $R_{AB}$ values suggests differences in baryon versus meson production instead of a simple mass dependence. In $p$$+$Al collisions proton $R_{AB}$ values and $R_{AB}$ values of all measured mesons are in agreement within uncertainties, which shows zero enhancement in proton to $\phi$-meson production.
@@ IV SUMMARY
The PHENIX experiment has measured identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra and nuclear-modification factors $R_{AB}$, and $p/\pi$, $K/\pi$ ratios in $p$$+$Al, ${}^{3}$e$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV.
The values of freeze-out temperatures $T_{0}$ and average collective velocities $u_{t}\right>$ have been obtained. The $T_{0}$ values do not exhibit any dependence on the collision centrality and $\langle N_{\rm part}\rangle$ values, whereas the values of $u_{t}\right>$ smoothly increase with increasing of $\langle N_{\rm part}\rangle$ values. This indicates that in collisions characterized by large $\langle N_{\rm part}\rangle$ values (central Cu$+$Au, Au$+$Au, U$+$U collisions), collective effects are more pronounced than in collision systems with small $\langle N_{\rm part}\rangle$ values ($p$$+$Al, ${}^{3}$e$+$Au collisions and peripheral collisions of large systems).
@@
In heavy ion collisions (Cu$+$Au and U$+$U) the $p$/$\pi$ ratios exhibit strong centrality dependence, but in small ${}^{3}$e$+$Au collisions the centrality dependence is much more modest due to the much smaller range of $\langle N_{\rm part}\rangle$ values. Within uncertainties, the $p$/$\pi$ values measured in each centrality of $p$$+$Al collisions are consistent with those measured in $p$$+$$p$ collisions. The values of the $K/\pi$ ratios show a modest centrality dependence, which is insignificant within systematic uncertainties. The observed behavior of $p/\pi$ and $K/\pi$ ratios can be qualitatively described through hadronization by recombination [23, 24].
Comparison of identified charged-hadron nuclear-modification factors shows that in $d$$+$Au, ${}^{3}$e$+$Au, Cu$+$Au, Au$+$Au, and U$+$U collision systems the $R_{AB}$ values are consistent at the same number of participant nucleons $\langle N_{\rm part}\rangle$. This indicates that identified charged-hadron production does not depend on the geometry and collision species, but rather is determined by system size alone (as indicated by $\langle N_{\rm part}\rangle$ values). Further, it is found that: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than in ${}^{3}$e$+$Au and $d$$+$Au collisions at the same $\langle N_{\rm part}\rangle$ values, (ii) proton $R_{AB}$ values in $p$$+$Al collisions are equal to unity in the range of 1.0 GeV/$c$ $<p_{T}<$ 2.5 GeV/$c$, while proton $R_{AB}$ values measured in ${}^{3}$e$+$Au and $d$$+$Au collisions are larger than unity. Despite the observed absence of proton enhancement in $p$$+$Al collisions, there were evidences of QGP formation found by the PHENIX experiment in studies of $J/\psi$, $\psi(2S)$, and charged-hadron production at backward rapidity [48]. The observed differences in identified charged hadron production in $p$$+$Al and $d$/${}^{3}$e$+$Au collisions may be caused by the size of the system created in $p$$+$Al collisions being too small to observe the expected increase in proton production.
@@ ACKNOWLEDGMENTS
We thank the staff of the Collider-Accelerator and Physics Departments at Brookhaven National Laboratory and the staff of the other PHENIX participating institutions for their vital contributions. We acknowledge support from the Office of Nuclear Physics in the Office of Science of the Department of Energy, the National Science Foundation, Abilene Christian University Research Council, Research Foundation of SUNY, and Dean of the College of Arts and Sciences, Vanderbilt University (U.S.A), Ministry of Education, Culture, Sports, Science, and Technology and the Japan Society for the Promotion of Science (Japan), Natural Science Foundation of China (People’s Republic of China), Croatian Science Foundation and Ministry of Science and Education (Croatia), Ministry of Education, Youth and Sports (Czech Republic), Centre National de la Recherche Scientifique, Commissariat à l’Énergie Atomique, and Institut National de Physique Nucléaire et de Physique des Particules (France), J. Bolyai Research Scholarship, EFOP, the New National Excellence Program (ÚNKP), NKFIH, and OTKA (Hungary), Department of Atomic Energy and Department of Science and Technology (India), Israel Science Foundation (Israel), Basic Science Research and SRC(CENuM) Programs through NRF funded by the Ministry of Education and the Ministry of Science and ICT (Korea). Ministry of Education and Science, Russian Academy of Sciences, Federal Agency of Atomic Energy (Russia), VR and Wallenberg Foundation (Sweden), University of Zambia, the Government of the Republic of Zambia (Zambia), the U.S. Civilian Research and Development Foundation for the Independent States of the Former Soviet Union, the Hungarian American Enterprise Scholarship Fund, the US-Hungarian Fulbright Foundation, and the US-Israel Binational Science Foundation.

### v2

@@ III.1 Particle ratios
Enhancement of baryon production in nucleus-nucleus collisions is considered to be one of the signatures of QGP formation [14, 15, 17]. To investigate differences in baryon and meson production mechanisms, the ratios of $p/\pi$ and $K/\pi$ have been calculated. Figures 9 and 9 present comparisons of $p$/$\pi^{+}$ and $\bar{p}$/$\pi^{-}$ ratios in different centralities of large (Cu$+$Au, U$+$U) and small ($p$$+$$p$, $p$$+$Al, 3He$+$Au) collision systems. In central collisions of large systems $p$/$\pi$ ratios reach the values of  $\approx\,0.6$, but in peripheral collisions the values of $p/\pi$ ratios are smaller than 0.4 in the whole $p_{T}$ range. Behavior of $p/\pi$ ratios observed in Cu$+$Au and U$+$U collision systems can be qualitatively described using recombination models [23, 24].
In small collision systems ($p$$+$Al, 3He$+$Au), the values of $p$/$\pi$ ratios are similar to those measured in $p$$+$$p$ collisions [43]. In 3He$+$Au collisions a modest centrality dependence can be seen, similar to that observed in $d$$+$Au collisions [14, 17]. The modest centrality dependence can be understood in terms of the small range of $\langle N_{\rm part}\rangle$ values relative the large range of $\langle N_{\rm part}\rangle$ values in large collision systems. The values of $p/\pi$ ratios measured in all centrality classes of $p$$+$Al collisions and in $p$$+$$p$ collisions are consistent within uncertainties. For all measurements type A uncertainties sum quadratically to type B uncertainties and are shown as a rectangles around the experimental points in Figs. 1115.
@@ III.2 Nuclear-modification factors
Figures 13 and 13 present comparisons of identified charged-hadron $R_{AB}$ values as a function of $p_{T}$ in central and peripheral $p$$+$Al, 3He$+$Au, Cu$+$Au, and U$+$U collisions. The $R_{AB}$ values are found to be in agreement in collisions with different geometries, but with the same $\langle N_{\rm part}\rangle$ values, indicating that identified charged-hadron production depends only on system size and not geometry.
The following features of identified charged-hadron production in $p$$+$Al collisions have been found: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than it is in 3He$+$Au and $d$$+$Au collisions and (ii) proton $R_{AB}$ values in $p$$+$Al collisions at the intermediate $p_{T}$ range (1.0 GeV/$c$ $<~{}p_{T}~{}<$ 2.5 GeV/$c$) are equal to unity, while in 3He$+$Au and $d$$+$Au collisions proton $R_{AB}$ is above unity. Differences in identified charged-hadron production between $p$$+$Al and $d/$3He$+$Au might be caused by the size of the $p$$+$Al system being insufficient to observe an increase in proton production.
Comparisons of identified charged-hadron $R_{AB}$ values with neutral-meson $R_{AB}$ values [47, 44, 45] in small and large collision systems are presented in Figs. 15 and 15. In large collision systems and in the 3He+Au collision system, proton $R_{AB}$ values are enhanced over all meson $R_{AB}$ values. The mass of the $\phi$-meson $m_{\phi}$=1019 MeV/$c^{2}$ is similar to the proton mass $m_{p}$=938 MeV/$c^{2}$, therefore the enhancement of proton $R_{AB}$ values over $\phi$-meson $R_{AB}$ values suggests differences in baryon versus meson production instead of a simple mass dependence. In $p$$+$Al collisions proton $R_{AB}$ values and $R_{AB}$ values of all measured mesons are in agreement within uncertainties, which shows zero enhancement in proton to $\phi$-meson production.
Figure 16: The $\pi^{+}$, $\pi^{-}$, $K^{+}$,$K^{-}$, $p$ and $\bar{p}$ invariant $p_{T}$ spectra measured in different centralities of $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV. Invariant $p_{T}$ spectra are multiplied by powers of ten for clarity of presentation.
The $p/\pi^{+}$ and $\bar{p}/\pi^{-}$ ratios integrated in the low-$p_{T}$ (high-$p_{T}$) regions, $p_{T}<1.5$ GeV/$c$ ($p_{T}>1.5$ GeV/$c$) are plotted in Fig. 16 as a function of $N_{\rm part}\right>$. The values of ratios, integrated in the low-$p_{T}$ region [Fig. 16(a) and 16(c)], were found to be approximately independent of $N_{\rm part}\right>$ values. The values of ratios, integrated in the high-$p_{T}$ region [Fig. 16(b) and 16(d)], smoothly grow with increasing $N_{\rm part}\right>$. Because the $\bar{p}/p$ ratio is approximately equal to 0.73 [14], regardless of $N_{\rm part}\right>$ the integrated $p/\pi^{+}$ ratios exceed the integrated $\bar{p}/\pi^{-}$ ratios.
According to the recombination model, the $p_{T}$ of a produced hadron is the sum of the $p_{T}$ of its constituent quarks. Therefore, baryon-invariant $p_{T}$ spectra are shifted towards larger $p_{T}$ relative to the meson $p_{T}$ spectra, which leads to an enhancement in baryon over meson production in the intermediate $p_{T}$ region. Therefore, the behavior of the integrated ratio supports the assumption that the influence of the recombination mechanism grows with increasing $N_{\rm part}\right>$.
@@ IV SUMMARY
The PHENIX experiment has measured identified charged-hadron invariant $p_{T}$ and $m_{T}$ spectra and nuclear-modification factors $R_{AB}$, and $p/\pi$, $K/\pi$ ratios in $p$$+$Al, 3He$+$Au, Cu$+$Au collisions at $\sqrt{s_{{}_{NN}}}$ = 200 GeV and in U$+$U collisions at $\sqrt{s_{{}_{NN}}}$ = 193 GeV.
The values of freeze-out temperatures $T_{0}$ and average collective velocities $u_{t}\right>$ have been obtained. The $T_{0}$ values do not exhibit any dependence on the collision centrality and $\langle N_{\rm part}\rangle$ values, whereas the values of $u_{t}\right>$ smoothly increase with increasing of $\langle N_{\rm part}\rangle$ values. This indicates that in collisions characterized by large $\langle N_{\rm part}\rangle$ values (central Cu$+$Au, Au$+$Au, U$+$U collisions), collective effects are more pronounced than in collision systems with small $\langle N_{\rm part}\rangle$ values ($p$$+$Al, 3He$+$Au collisions and peripheral collisions of large systems).
@@
In heavy ion collisions (Cu$+$Au and U$+$U) the $p$/$\pi$ ratios exhibit strong centrality dependence, but in small 3He$+$Au collisions the centrality dependence is much more modest due to the much smaller range of $\langle N_{\rm part}\rangle$ values. Within uncertainties, the $p$/$\pi$ values measured in each centrality of $p$$+$Al collisions are consistent with those measured in $p$$+$$p$ collisions. The values of the $K/\pi$ ratios show a modest centrality dependence, which is insignificant within systematic uncertainties. The observed behavior of $p/\pi$ and $K/\pi$ ratios can be qualitatively described through hadronization by recombination [23, 24].
Comparison of identified charged-hadron nuclear-modification factors shows that in $d$$+$Au, 3He$+$Au, Cu$+$Au, Au$+$Au, and U$+$U collision systems the $R_{AB}$ values are consistent at the same number of participant nucleons $\langle N_{\rm part}\rangle$. This indicates that identified charged-hadron production does not depend on the geometry and collision species, but rather is determined by system size alone (as indicated by $\langle N_{\rm part}\rangle$ values). Further, it is found that: (i) the slope of $R_{AB}(p_{T})$ in $p$$+$Al collisions is flatter than in 3He$+$Au and $d$$+$Au collisions at the same $\langle N_{\rm part}\rangle$ values, (ii) proton $R_{AB}$ values in $p$$+$Al collisions are equal to unity in the range of 1.0 GeV/$c$ $<p_{T}<$ 2.5 GeV/$c$, while proton $R_{AB}$ values measured in 3He$+$Au and $d$$+$Au collisions are larger than unity. Despite the observed absence of proton enhancement in $p$$+$Al collisions, there were evidences of QGP formation found by the PHENIX experiment in studies of $J/\psi$, $\psi(2S)$, and charged-hadron production at backward rapidity [48]. The observed differences in identified charged hadron production in $p$$+$Al and $d$/3He$+$Au collisions may be caused by the size of the system created in $p$$+$Al collisions being too small to observe the expected increase in proton production.
@@ ACKNOWLEDGMENTS
We thank the staff of the Collider-Accelerator and Physics Departments at Brookhaven National Laboratory and the staff of the other PHENIX participating institutions for their vital contributions. We acknowledge support from the Office of Nuclear Physics in the Office of Science of the Department of Energy, the National Science Foundation, Abilene Christian University Research Council, Research Foundation of SUNY, and Dean of the College of Arts and Sciences, Vanderbilt University (U.S.A), Ministry of Education, Culture, Sports, Science, and Technology and the Japan Society for the Promotion of Science (Japan), Natural Science Foundation of China (People’s Republic of China), Croatian Science Foundation and Ministry of Science and Education (Croatia), Ministry of Education, Youth and Sports (Czech Republic), Centre National de la Recherche Scientifique, Commissariat à l’Énergie Atomique, and Institut National de Physique Nucléaire et de Physique des Particules (France), J. Bolyai Research Scholarship, EFOP, HUN-REN ATOMKI, NKFIH, and OTKA (Hungary), Department of Atomic Energy and Department of Science and Technology (India), Israel Science Foundation (Israel), Basic Science Research and SRC(CENuM) Programs through NRF funded by the Ministry of Education and the Ministry of Science and ICT (Korea). Ministry of Education and Science, Russian Academy of Sciences, Federal Agency of Atomic Energy (Russia), VR and Wallenberg Foundation (Sweden), University of Zambia, the Government of the Republic of Zambia (Zambia), the U.S. Civilian Research and Development Foundation for the Independent States of the Former Soviet Union, the Hungarian American Enterprise Scholarship Fund, the US-Hungarian Fulbright Foundation, and the US-Israel Binational Science Foundation.